OpenStudy (surfkitty05):

HELP Please! Verify the identity. (Show work) Cot 0 (x) sec (x) = csc =0

OpenStudy (error1603):

(sin x cot x)/(cos x csc x) = (sin x • cos x/sin x)/(cos x • 1/sin x) = cos x • sin x/cos x = sin x (1 + csc x)(sec x - tan x) = (1 + 1/sin x)(1/cos x - sin x/cos x) = (sin x + 1)/sin x • (1 - sin x)/cos x = (1 - sin^2 x)/(sin x cos x) = cos^2 x/(sin x cos x) = cos x/sin x = cot x

OpenStudy (misty1212):

HI!

OpenStudy (surfkitty05):

the (x) is actually a multiplication dot

OpenStudy (misty1212):

cotangent is cosine over sine secant is one over cosine

OpenStudy (misty1212):

\[\frac{\cos(\theta)}{\sin(\theta)}\times \frac{1}{\cos(\theta)}\]cancel the cosines

OpenStudy (surfkitty05):

ohh, thank you :)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]